Energy consumed in the US can be classified as

coming from one of three sources: fossil fuels, nuclear power, and

renewable energy. In 2014, the energy from these three sources

was 80.3, 8.3, and 9.6 quadrillion BTU, respectively. In 2004, the

corresponding amounts were 85.8, 8.2, and 6.1. Write a description

of the changes from 2004 to 2014 expressed in these data. Illustrate

your summary with appropriate graphical summaries. Be sure to

discuss both the amounts of energy from each source as well as the

percents.
Answer #1

The data values for each energy sources are given as shown below,

Energy source 2004 2014
fossil fuels 80.3 85.8
nuclear power 8.3 8.2
renewable energy 9.6 6.1

Graphical display of changes from 2004 to 2014

To see the changes in the graph, the bar plot is obtained in excel by following these steps,

Step 1: Write the data values in excel. The screenshot is shown below,

Step 2: Select all the data values and go to chart area then INSERT > Recommended Chart > Clustered Bar > OK. The screenshot of the chart is shown below,

Conclusion: From the bar chart we can observe that the generation of the amount of fossil fuel is decreased, the amount of renewable energy is increased while the amount of nuclear power is approximately the same from the year 2004 to 2014.

Percent change in the energy source (with respect to previous value) from 2004 to 2014

The % change in the amount of energy source is obtained using the following formula,

Energy source 2014 2004 % Change
fossil fuels 80.3 85.8 - 6.85
nuclear power 8.3 8.2 1.20
renewable energy 9.6 6.1 36.46

Conclusion: From the above table we can observe that the source of renewable energy is increased by 36.46%. from the previous value, source of fossil fuel is decreased by 6.85%. from the previous value and the source of nuclear energy is increased by 1.20%. from the previous value.

Overall percent change in the energy source (with respect to the total value of energy source) from 2004 to 2014

The % share of each energy source for the year 2004 is obtained by dividing the amount of the energy source by the total energy source.

Energy source 2014 % Share 2004 % Share Change in the % share
fossil fuels 80.3 80.3*100/98.2=81.77 85.8 85.8*100/100.1=85.71 -3.94
nuclear power 8.3 8.3*100/98.2=8.45 8.2 8.2*100/100.1=8.19 0.26
renewable energy 9.6 9.6*100/98.2=9.78 6.1 6.1*100/100.1=6.09 3.68
Total 98.2 100.1

The overall share of renewable energy is increased by 3.68%, the overall share of fossil fuel is decreased by 3.94%, and the overall share of nuclear power is increased by 0.26%.